OpenStudy (anonymous):

MEDAL! what is 1/2 *12*8? please explain how to do this

OpenStudy (anonymous):

@sleepyjess

OpenStudy (sleepyjess):

What is 12 as a fraction?

OpenStudy (anonymous):

no, I know that the equation should look like this----> 1/2* 12/1 * 8/1 but I don't know how to solve it

OpenStudy (anonymous):

that doesn't work in this case @K_V8

OpenStudy (sleepyjess):

Ok, that is correct. \(\dfrac{1}{2}*\dfrac{12}{1}*\dfrac{8}{1}\). First, lets just focus on \(\dfrac{1}{2}*\dfrac{12}{1}\). Mulitply numerator*numerator, then denominator*denominator.

OpenStudy (anonymous):

hmmm.

OpenStudy (anonymous):

it equals.... 6?

OpenStudy (anonymous):

@sleepyjess

OpenStudy (sleepyjess):

Yep, \(\dfrac{12}{2}=\dfrac{6}1\). Now, \(\dfrac{6}{1}*\dfrac{8}{1}\)

OpenStudy (anonymous):

48/1?

OpenStudy (anonymous):

@sleepyjess

OpenStudy (sleepyjess):

Yep, which is equal to 48 :)

OpenStudy (anonymous):

its 48/1

OpenStudy (anonymous):

thank you! I thought you had to do this----> 1/2 * 12/1=12/2=6 and multiply the same way for 8/1 which would be 4 and I did 6*4= 24 and I thought that was the answer

OpenStudy (sleepyjess):

@Avery13 Welcome to OpenStudy! Please do not give out direct answers.

OpenStudy (sleepyjess):

No problem @kittenlover731 :)